OpenStudy (anonymous):

Find all relative extrema, use the second derivative test where applicable... f(x)=6x-x^2

OpenStudy (freckles):

well have you found f'?

OpenStudy (freckles):

once we find f' we should find for what x values f'=0 ( i would say f' dne but we don't have to work about that case) then once we find f' we can find derivative of f' to get f''

OpenStudy (anonymous):

yeah i already got all of that, f''(x)=-2

OpenStudy (anonymous):

So from there where do i go..

OpenStudy (anonymous):

and the critical point is x=3

OpenStudy (freckles):

so f'=6-2x f'=0 when x=3 and f''=-2 f''=-2 means your graph is open down x=3 so what does imply what happens at x=3?

OpenStudy (anonymous):

right, its a relative max @ 3

OpenStudy (freckles):

 so what i'm saying is your picture looks sorta like this and yes